OpenStudy (anonymous):

. Choose the correct alternative that will continue the same pattern and fill in the blank spaces. 2, 7, 14, 23, ?, 47 A. 31 B. 28 C. 34 D. 38

OpenStudy (anonymous):

check the difference forms an arithmetic sequence.. 7-2=5 14-7=7 23-14=9 ?-23 =11 so the reqd number is 34

OpenStudy (lncognlto):

It goes +5, +7, +9, so the next is +11. So the answer is C, 34.

OpenStudy (anonymous):

4, 6, 12, 14, 28, 30, ? A. 32 B. 64 C. 62 D. 60 Here the answer would be 60? Right?

OpenStudy (lncognlto):

6 - 4 = 2 12/6 = 2 14 - 12 = 2 28/14 = 2 So, yes, 30 x 2 = 60, which is the right answer.

OpenStudy (anonymous):

I need some help with some coding-decoding questions.. Can you explain me what I need to do? 1) If FRIEND is coded as HUMJTK, how can CANDLE be written in that code? A. DEQJQM B. DCQHQK C. EDRIRL D. ESJFME

OpenStudy (lncognlto):

Okay, here is how it works. Writing out the alphabet, ABCDEFGHIJKLMNOPQRSTUVWXYZ, and then noting how each letter in FRIEND changes to HUMJTK: F +2, R +3, I +4, E +5, N + 6, and D +7. So then CANDLE would be C +2, A +3, N +4, D +5, L +6, and E +7.

OpenStudy (anonymous):